# Optimal Transport Aggregation for Visual Place Recognition

### v1

@@ 4.4 Introspective Results
We provide an introspection of our model’s performance through a series of illustrative figures. Figure 4 visualizes the weights that are not assigned to the ‘dustbin’. This visualization offers insight into the parts of the input image that the network considers informative and validates the effective use of the ‘dustbin’ by our SALAD aggregation. In Figure 6, we display the assignment distribution of patches from two different images depicting the same place. It demonstrates the model’s ability to consistently distribute most of the weights into the same bins for patches representing similar regions. Such repeatable and consistent assignment across different images of the same place is crucial for the reliability and performance of the system. Finally, in Figure 6, we showcase various query images alongside their respective top-3 retrievals made by our system. DINOv2 SALAD is able to retrieve correct predictions even under challenging conditions, such as severe changes in illumination or viewpoint.
@@
Figure 6: DINOv2 SALAD qualitative results at MSLS. The left column shows several queries and the three other ones shows the top-3 candidates retrieved by our DINOv2 SALAD. Candidates are framed in green if they correspond to the same place as the query, and in red if they do not. Note the correct retrievals under seasonal, weather, viewpoint and day-night changes. Note also a challenging failure case in the last row, due to non-disciminative image content.
@@ 5 Conclusions and Limitations
Our research primarily focused on standard benchmarks, predominantly outdoor environments. The use of the DINOv2 backbone, while effective in these contexts, might encounter limitations in vastly different scenarios from DINOv2’s training distribution, such as medical images. Additionally, in SALAD we use an optimal transport assignment in its simplest form. More sophisticated constraints could improve the resulting assignment, a very relevant aspect for our future work.

### v2

@@ 4.4 Introspective Results
We provide an introspection of our model’s performance through a series of illustrative figures. Figure 3 visualizes the weights that are not assigned to the ‘dustbin’, offering insight into the parts of the input image that the network considers informative. As the ‘dustbin’ assignment is completely learnt by the network, some discarded features might be counter-intuitive. However, we observe that it typically removes dynamic objects and focuses on the most distinctive and invariant parts of the image. In Figure 5, we display the assignment distribution of patches from two different images depicting the same place. It demonstrates the model’s ability to consistently distribute most of the weights into the same bins for patches representing similar regions. Such repeatable and consistent assignment across different images of the same place is crucial for the reliability and performance of the system. Finally, in Figure 5, we showcase various query images alongside their respective top-3 retrievals made by our system. DINOv2 SALAD is able to retrieve correct predictions even under challenging conditions, such as severe changes in illumination or viewpoint.
@@
Figure 5: DINOv2 SALAD qualitative results at MSLS. The left column shows several queries and the three other ones shows the top-3 candidates retrieved by our DINOv2 SALAD. Candidates are framed in green if they correspond to the same place as the query, and in red if they do not. Note the correct retrievals under seasonal, weather, viewpoint and day-night changes. Note also a challenging failure case in the last row, due to non-discriminative image content.
@@ 5 Conclusions and Limitations
While our work brings significant improvements in performance, it is not without limitations. Primarily, the adoption of DINOv2 as our backbone results in slower processing speeds compared to ResNet-based methods. Besides, although SALAD is a general aggregation module, its effectiveness is tied to the choice of backbone. It excels with DINOv2, which offers high spatial resolution features, but it is less suited for coarser features. Additionally, in SALAD we use an optimal transport assignment in its simplest form. More sophisticated constraints could improve the resulting assignment, a very relevant aspect for our future work.
Acknowledgments
This work was supported by the Spanish Government (PID2021-127685NB-I00 and TED2021-131150B-I00), the Aragón Government (DGA T45_23R) and the scholarship FPU20/02342.